# Quantum simulating continuum field theories with large-spin lattice models

### v1

@@ II.2.3 Matching microscopic and QFT parameters
which provides the required direct relation of the bare dimensionless parameters $\beta^{\prime}$ and $M_{0}^{\prime}=aM_{0}$ of the QFT to the microscopic parameters of the lattice model [see App. B for more details]. Note that this identification requires tuning
@@ II.3 Accessible QFTs
$V(\hat{\varphi})=-\frac{M_{0}^{2}}{\beta^{2}}\sum_{\tilde{\kappa}}\frac{\lambda^{\prime}_{\tilde{\kappa}}}{\lambda^{\prime}_{\kappa}}\cos\left(\beta\frac{\tilde{\kappa}}{\kappa}\hat{\varphi}\right)\;,$ (11)
@@
For larger spatial dimensions ($d>1$), scalar fields have a non-vanishing mass dimension, namely $[\hat{\varphi}]=\frac{d-1}{2}$ in units of $[M_{0}]=1$. According to textbook arguments [49], renormalizable field theories in $d=2$ and $d=3$ include interaction potentials up to $\sim\hat{\varphi}^{6}$ and $\sim\hat{\varphi}^{4}$, respectively. These define two interesting target models, where quantum simulators could tackle long-standing questions in non-equilibrium QFT about the fate of the false vacuum [50]. For recent alternative approaches to this problem, we refer to the literature [51, 52, 53, 43].
@@ II.4 Overview of results
In Sec. III, we show with detailed numerical tensor-network simulations that our proposed large-spin regularization faithfully captures equilibrium properties of the 1D sine-Gordon model in the continuum limit. Focusing first on the underlying conformal field theory (CFT), the Luttinger liquid model, we simulate several system sizes $N$ and spin lengths $J$, and perform the sequence of extrapolations discussed in Sec. II.2 [see also Fig. 2], to identify the parameter regime of the predicted gapless phase with associated central charge $c=1$. We further determine the Luttinger parameter $K$ of the field theory and clarify its renormalization w.r.t. the classical microscopic mapping from the lattice Hamiltonian.
@@
We then consider the sine-Gordon model where we first determine the values of the renormalized coupling strengths $\beta^{2},M_{0}^{\prime}$, confirming the theoretical prediction $\beta^{2}=\pi\kappa^{2}K$ between renormalized parameters of the sG and the CFT. By performing the sequence of limits in $J$ and $N$, we extract the scaling behaviour of several vacuum and first-excited state observables in the continuum limit, including the mass gap and vertex operators, in the regime $\beta^{2}\lesssim 1$. Our numerical results demonstrate quantitative agreement with theoretical QFT predictions for the renormalized coupling parameters, and also allows us to test the unproven Lukyanov-Zamolodchikov conjecture. Overall, we find that the investigated system sizes of $N\lesssim 100$ lattice sites and spin lengths $J\lesssim 20$ – both within reach of state-of-the-art quantum simulators – are adequate for the equilibrium quantum simulation of the sG model. These results form the basis for the following sections where we apply our approach to the investigation of localized excitations, their dynamics and collisions in the sG model.
@@
In Sec. IV, we simulate the real-time quantum dynamics of individual solitons for small $\beta^{2}$. We propose a simple preparation scheme of “semi-classical solitons”, i.e., quantum states whose phase profile agrees with the classical soliton solution, but are subject to the same fluctuations dictated by quantum correlations in the true ground states. Assuming local control over the microscopic spins, this preparation protocol applies local spin rotations on the ground state of our lattice regularization of the quantum sG model to adjust the one-point field expectation values while preserving the two-point functions. Starting from such initial states, we find that the time evolution displays the expected classical linear motion in the mean position of the soliton, with a velocity that agrees with the one imprinted by our protocol. On top of this classical motion, the presence of quantum fluctuations induces other observables – such as the topological charge density – to spread and smear out over time. This observation can be explained by interpreting the prepared states as wave-packets of solitons, whose initial velocities and positions are stochastically distributed around the imposed classical values. We make this interpretation precise by constructing an appropriate semi-classical phenomenological model, essentially a truncated Wigner approximation restricted to the single-soliton sector, from which we derive analytical predictions that quantitatively agree with the observed dynamics.
@@
In Sec. V, we further leverage the state preparation tools developed in the previous section and simulate the scattering of a soliton and antisoliton. Within the parameter regimes accessible to our classical numerical simulations, we observe how the two excitations initially counter-propagate, with the expected linear motion, followed by transmissive scattering, after which the individual excitations continue to propagate linearly. The considered type of collision is expected to introduce a position shift of the outgoing trajectories compared to the incoming ones, which we extract from our simulation. Our results agree with predictions from the known S-matrix of the sG model for several different initial conditions, verifying that also the non-equilibrium dynamics is faithfully captured by the large-spin regularization.
@@
In Sec. VI, we consider a variant of our proposed regularization that corresponds to a non-integrable perturbation of the sG model. The form of the perturbation is motivated by quantum electrodynamics and is expected to introduce a confining potential between soliton and antisoliton, which in turn leads to the formation of a meson-like bound state. We classically simulate scattering dynamics of this modified model, which reveals qualitatively distinct regimes depending on the perturbation strength. Small integrability breaking leads to repeated scattering of the individual quasi-particles, consistent with the formation and oscillations of a bound state. Upon increasing the perturbation strength, we observe dynamics reminiscent of string breaking, as suggested by the pairwise production of quasi-particles of opposite topological charge, as well as “plasma” oscillation of observables akin to the electric field strength. While a quantitative analysis of this dynamics at long times is prohibitively costly with classical means, this setup is a prime example for the application of our approach in a quantum simulation experiment.
@@ III.1 The sG model in the continuum
[…]
@@
[…]
@@ III.2 The sG model from the lattice
We now turn to numerical simulations of the lattice regularization with finite-length spins. Throughout this work, we employ matrix product states (MPS) techniques using ITensor [63]. In order to obtain continuum results, we need to carry out the extrapolations discussed in the previous section. In practice, we do this by working with a fixed $\chi=1$ and scanning several values of system size $N=30,\dots,100$, spin length $J=12,\dots,20$ and couplings $V_{\text{nn}},\lambda_{\kappa}$. Overall, we find our simulations to be well described by first fitting the numerical data with a linear function in $1/J(J+1)$, and then with a quadratic function in $1/N$, see Fig. 2 and the discussion in App. C. Unless stated otherwise, we present observables obtained after this two-step extrapolation ($J\rightarrow\infty$ followed by $N\rightarrow\infty$) and only highlight the more delicate continuum scaling from now on.
@@ III.2.1 The underlying CFT
$0<\frac{\chi}{V^{\prime}_{\text{nn}}}<\left(\frac{\chi}{V^{\prime}_{\text{nn}}}\right)_{\text{crit}}\approx 0.27$ (16)
@@
To test the nature of the anticipated CFT at small $\chi/V^{\prime}_{\text{nn}}$, we further calculate the von Neumann entanglement entropy $S_{\text{vN}}$ in the ground state for a bipartition of $n$ sites and their complement of $N-n$ sites and fit the central charge $c$ according to
@@
with $c^{\prime}$ a non-universal constant [64] [see App. D]. Our results confirm $c=1$ as expected for a Luttinger liquid [see inset of Fig. 3(a)].
@@
These expectations are also confirmed by our simulations as shown in Fig. 3(b), at least for sufficiently small $K$. For larger $\sqrt{\chi/V^{\prime}_{\text{nn}}}\gtrsim 0.4$, we find quantitative deviations, which we attribute to higher derivative terms appearing in the expansion of $\cos\left[\beta(\hat{\varphi}_{i}-\hat{\varphi}_{j})/\kappa\right]=1-\beta^{2}(\hat{\varphi}_{i}-\hat{\varphi}_{j})^{2}/(2\kappa^{2})+\dots$ from the dipole-dipole interaction $\hat{J}_{+}^{(i)}\hat{J}_{-}^{(j)}+\mathrm{H.c.}\;$. Our numerics thus shows that these neglected higher-order terms can be absorbed in a renormalization of the Luttinger parameter $K$. For quantitative predictions in a potential experimental realization, one could similarly extract $K$ (which determines the coupling $\beta$ in the sG model through $\beta^{2}=\pi\kappa^{2}K$) by a fit of Eq. (III.2.1) instead of relying on the above theoretical estimates. We also emphasize that larger values of $\beta^{2}$ become accessible by realizing interactions with $\kappa>1$.
@@ III.2.2 The sine-Gordon model
which we test in our numerics by a linear fit in log-log scale. The result is shown in Fig. 4(a), where we compare the fit results with the scaling dimensions extracted from the CFT. The observed agreement corroborates our identification of the renormalized coupling $\beta^{2}$.
Since the sG model has no further coupling parameters, we are left with the renormalization scale, which is set by the bare mass $M_{0}^{\prime}$ on the microscopic lattice. Given the observed renormalization of $\beta^{2}$ due to higher derivatives from the term $\propto\cos\left[\beta\left(\hat{\varphi}_{i}-\hat{\varphi}_{j}\right)/\kappa\right]$ [see App. E.2], we depart from the microscopic identification given in Eq. (9). Instead, we fix $M_{0}^{\prime}$ through the formally equivalent relation
@@
which is, however, independent of the microscopic parameter $V^{\prime}_{\text{nn}}$. We obtain an adjusted prediction of $M_{0}^{\prime}$ by inserting the value of $\beta^{\prime}=\beta$ from the previously extracted $\beta^{2}$ [either through $\beta^{2}=\pi\kappa^{2}K$ or Eq. (20)]. As shown in Fig. 2(c), the numerically obtained mass gap $\Delta E$ as a function of mass scale set by this procedure shows excellent agreement with the analytical predictions from Eqs. (12) and (13). This clearly indicates that our numerical results have reached the scaling limit and indeed realize the sG QFT with correctly extracted parameters $M_{0}^{\prime}$ and $\beta^{2}$. Further details about the matching of the mass scale can be found in App. E.3.
Figure 4: Renormalized sG QFT. (a) Coupling $\beta^{2}$ as a function of the microscopic parameters $\kappa^{2}\sqrt{\chi/V_{\text{nn}}^{\prime}}$ extracted in the sG model for $\kappa=1$ (colored dots) and $\kappa=2$ (yellow diamonds) with a fit according to Eq. (20) [see inset]. The extracted coupling parameters agree with the prediction from the CFT according to $\beta^{2}=\pi\kappa^{2}K$ (gray squares), but show deviations from the naive estimate from the microscopic parameters [Eq. (9), dashed black line], signalling a renormalization of $\beta^{2}$. (b) Ground-state energy density $E_{0,\text{latt}}$ (in lattice units) measured with respect to the CFT. For the same value of $\beta^{2}$ we display the energy gap $\Delta E_{\text{sG}}$ (in sG units) in the inset. (c) Expectation value of the vertex operator $\langle e^{i\beta\hat{\varphi}}\rangle$ at $\kappa=1$ versus $M_{0}^{\prime}$. Numerical results (colored plus) and corresponding theoretical predictions (dashed black line). As discussed in the main text, the bare numerical result is renormalized with a factor $Z_{\beta}\approx 1+\mathcal{O}(\beta^{2})$ [see inset]. (d) The expectation value of higher powers of the vertex operators $\langle e^{in\beta\hat{\varphi}}\rangle$ (dark plus) at $\lambda^{\prime}_{\kappa}/V_{\text{nn}}^{\prime}=0.0025$ agree with the Lukyanov-Zamolodchikov conjecture (green circles; dashed lines are guides to the eye) upon rescaling the numerical results with the factor $(Z_{\beta})^{n^{2}}$ [see main text and App. E.4 for more details].
@@
So far, we have compared our numerics to exact results of the sG QFT. Although the model is integrable, not all quantities are known exactly. In particular, there is an unproven conjecture about the expectation values of $\langle e^{in\beta\hat{\varphi}}\rangle$ for generic $n$ [see App. E.4 for the explicit formula]. Here, we test this conjecture in our simulations for integer $n=1,\dots,6$ [see Fig. 4(d)]. We find quantitative agreement with the conjectured functional form when taking into account a multiplicative factor $Z_{\beta}^{(n)}=\left(Z_{\beta}\right)^{n^{2}}$, which relates the higher powers of the vertex operators to their corresponding lattice regularization $\propto\left(\hat{J}_{+}\right)^{n}$.
Altogether, our findings confirm the expected simulability of the sG model regularized with large spins on a finite lattice. While the accessible couplings for $\kappa=1$ are limited to the massive regime with $\beta^{2}\ll 4\pi$, the comparison with our results for $\kappa=2$ [see Fig. 4(a)] also illustrates that the accessible region can be systematically enlarged by increasing $\kappa$. Specifically, according to Eq. (9), for $\kappa=5$ we expect to reach the sG model with large $\beta^{2}$ up to $\beta^{2}\gtrsim 8\pi$. While we have not investigated this regime classically because the extrapolations for larger $\kappa\gg 1$ require simulations at increasingly large spin length $J$, an experimental implementation could nevertheless probe this regime.

### v2

@@ II.2.3 Matching microscopic and QFT parameters
which provides the required direct relation of the bare dimensionless parameters $\beta^{\prime}$ and $M_{0}^{\prime}=aM_{0}$ of the QFT to the microscopic parameters of the lattice model [see App. C for more details]. Note that this identification requires tuning
@@ II.3 Accessible QFTs
[…]
@@
For larger spatial dimensions ($d>1$), scalar fields have a non-vanishing mass dimension, namely $\hat{\varphi}]=\frac{d-1}{2}$ in units of $M_{0}]=1$. According to textbook arguments [65], renormalizable field theories in $d=2$ and $d=3$ include interaction potentials up to $\sim\hat{\varphi}^{6}$ and $\sim\hat{\varphi}^{4}$, respectively. These define two interesting target models, where quantum simulators could tackle long-standing questions in non-equilibrium QFT about the fate of the false vacuum [66]. For recent alternative approaches to this problem, we refer to the literature [67, 68, 69, 59].
@@ II.4 Overview of results
In Sec. III, we show with detailed numerical tensor-network simulations that our proposed large-spin regularization faithfully captures equilibrium properties of the 1D sine-Gordon model in the continuum limit. Focusing first on the underlying conformal field theory (CFT), the Luttinger liquid model, we simulate several system sizes NN and spin lengths $J$, and perform the sequence of extrapolations discussed in Sec. II.2 [see also Fig. 2], to identify the parameter regime of the predicted gapless phase with associated central charge $c=1$. We further determine the Luttinger parameter $K$ of the field theory and clarify its renormalization w.r.t. the classical microscopic mapping from the lattice Hamiltonian.
@@
We then consider the sine-Gordon model where we first determine the values of the renormalized coupling strengths $\beta^{2},M_{0}^{\prime}$, confirming the theoretical prediction $\beta^{2}=\pi\kappa^{2}K$ between renormalized parameters of the sG and the CFT. We emphasize that our parameter matching procedures and the identification of microscopic with continuum variables does not rely on the solvability of the sG QFT, but is also directly applicable to the quantum simulation of other target models, without requiring input from theory simulations. By performing the sequence of limits in $J$ and $N$, we extract the scaling behaviour of several vacuum and first-excited state observables in the continuum limit, including the mass gap and vertex operators, in the numerically accessible regime $\beta^{2}\lesssim 1$. Our numerical results demonstrate quantitative agreement with theoretical QFT predictions for the renormalized coupling parameters, which we further use to test the unproven Lukyanov-Zamolodchikov conjecture. Overall, we find that the investigated system sizes of $N\lesssim 100$ lattice sites and spin lengths $J\lesssim 20$ – both within reach of state-of-the-art quantum simulators – are adequate for the equilibrium quantum simulation of the sG model. These results form the basis for the following sections where we apply our approach to the investigation of localized excitations, their dynamics and collisions in the sG model.
@@
In Sec. IV, we simulate the real-time quantum dynamics of individual solitons for small $\beta^{2}$, as a warm-up for studying collisions of single-particle wave-packets. We propose a simple preparation scheme of “semi-classical solitons”, i.e., quantum states whose phase profile agrees with the classical soliton solution, but are subject to the same fluctuations dictated by quantum correlations in the true ground states. Assuming local control over the microscopic spins, our preparation protocol applies local spin rotations on the ground state of our lattice regularization of the quantum sG model to adjust the one-point field expectation values while preserving the two-point functions. Starting from such initial states, we find that the time evolution displays the expected classical linear motion in the mean position of the soliton, with a velocity that agrees with the one imprinted by our protocol. On top of this classical motion, the presence of quantum fluctuations induces other observables – such as the topological charge density – to spread and smear out over time. This observation can be explained by interpreting the prepared states as wave-packets of solitons, whose initial velocities and positions are stochastically distributed around the imposed classical values. We make this interpretation precise by constructing an appropriate semi-classical phenomenological model, essentially a truncated Wigner approximation restricted to the single-soliton sector, from which we derive analytical predictions that quantitatively agree with the observed dynamics.
@@
In Sec. V, we further leverage the state preparation tools developed in the previous section and simulate the scattering of a soliton and antisoliton. Within the parameter regimes accessible to our classical numerical simulations, we observe how the two excitations initially counter-propagate, with the expected linear motion, followed by transmissive scattering, after which the individual excitations continue to propagate linearly. The considered type of collision is expected to introduce a position shift of the outgoing trajectories compared to the incoming ones, which we extract from our simulation. Our results agree with predictions from the known S-matrix of the sG model for several different initial conditions, verifying that also the non-equilibrium dynamics is faithfully captured by the large-spin regularization. In contrast to most previous works, which have focused on qualitative aspects of scattering dynamics, here we demonstrate quantitative agreement of real-time observables with continuum QFT results. A notable exception is the recent work Ref. [37], where the authors study real-time dynamics of scattering in a 1D Ising field theory, which is based on the well known Ising phase transition in a spin-1/2 chain. Similarly, Ref. [38] studies a realization of the sG model with two coupled spin-1/2 chains. In contrast, our approach using large spins enables the realization of general scalar fields in arbitrary dimensions.
@@
As an outlook illustrating the generality of our approach, we consider a variant of our proposed regularization that corresponds to a non-integrable perturbation of the sG model in Sec. VI. The form of the perturbation is motivated by quantum electrodynamics and is expected to introduce a confining potential between soliton and antisoliton, which in turn leads to the formation of a meson-like bound state. We classically simulate scattering dynamics of this modified model, which reveals qualitatively distinct regimes depending on the perturbation strength. Small integrability breaking leads to repeated scattering of the individual quasi-particles, consistent with the formation and oscillations of a bound state. Upon increasing the perturbation strength, we observe dynamics reminiscent of string breaking, as suggested by the pairwise production of quasi-particles of opposite topological charge, as well as “plasma” oscillation of observables akin to the electric field strength. While a quantitative analysis of this dynamics at long times is prohibitively costly with classical means, this setup is a prime example for the application of our approach in a quantum simulation experiment.
@@ III.1 The sG model in the continuum
The sine-Gordon (sG) model, defined by Eq. (8) for $d=1$, is a well-studied integrable relativistic 1+1D quantum field theory [47, 48, 49, 50, 51]. The dimensionless coupling $\beta^{2}=(\beta^{\prime})^{2}\geq 0$ identifies different regions of the model: at $\beta^{2}=8\pi$ a BKT transition separates a gapless phase ($\beta^{2}>8\pi$) from a gapped phase ($\beta^{2}\leavevmode\nobreak\ <\leavevmode\nobreak\ 8\pi$) [see Fig. 2(d)]. The gapped phase supports single-particle excitations known as (quantum) solitons whose mass $M$ is known analytically [50]
@@
where $\xi=\beta^{2}/(8\pi-\beta^{2})$, $\Lambda=\pi/a$ the UV cutoff, and $b=e^{C}/2$ a scale parameter dependent on the cutoff implementation. For the intermediate region $4\pi\leavevmode\nobreak\ <\leavevmode\nobreak\ \beta^{2}\leavevmode\nobreak\ <\leavevmode\nobreak\ 8\pi$ two solitons repel each other, while for $0\leavevmode\nobreak\ <\beta^{2}\leavevmode\nobreak\ <\leavevmode\nobreak\ 4\pi$ they interact attractively and form bound states (so-called breathers) of mass $m_{n}$ given by
@@ III.2 The sG model from the lattice
We now turn to numerical simulations of the lattice regularization with finite-length spins. Throughout this work, we employ matrix product states (MPS) techniques using ITensor [79]. In order to obtain continuum results, we need to carry out the extrapolations discussed in the previous section. In practice, we do this by working with a fixed $\chi=1$ and scanning several values of system size $N=30,\dots,100$, spin length $J=12,\dots,20$ and couplings $V_{\text{nn}},\lambda_{\kappa}$. Overall, we find our simulations to be well described by first fitting the numerical data with a linear function in $1/J(J+1)$, and then with a quadratic function in $1/N$, see Fig. 2 and the discussion in App. D. Unless stated otherwise, we present observables obtained after this two-step extrapolation ($J\rightarrow\infty$ followed by $N\rightarrow\infty$) and only highlight the more delicate continuum scaling from now on.
@@ III.2.1 The underlying CFT
$0<\frac{\chi}{V^{\prime}_{\text{nn}}}<\left(\frac{\chi}{V^{\prime}_{\text{nn}}}\right)_{\text{crit}}\approx 0.27.$ (16)
@@
To test the nature of the anticipated CFT at small $\chi/V^{\prime}_{\text{nn}}$, we further calculate the von Neumann entanglement entropy $S_{\text{vN}}$ in the ground state for a bipartition of $n$ sites and their complement of NnN-n sites and fit the central charge $c$ according to
@@
with $c^{\prime}$ a non-universal constant [80] [see App. E]. Our results confirm $c=1$ as expected for a Luttinger liquid [see inset of Fig. 3(a)].
@@
These expectations are also confirmed by our simulations as shown in Fig. 3(b), at least for sufficiently small $K$. For larger $\sqrt{\chi/V^{\prime}_{\text{nn}}}\gtrsim 0.4$, we find quantitative deviations, which we attribute to higher derivative terms appearing in the expansion of $\cos\left[\beta(\hat{\varphi}_{i}-\hat{\varphi}_{j})/\kappa\right]=1-\beta^{2}(\hat{\varphi}_{i}-\hat{\varphi}_{j})^{2}/(2\kappa^{2})+\dots$ from the dipole-dipole interaction $\hat{J}_{+}^{(i)}\hat{J}_{-}^{(j)}+\mathrm{H.c.}\;$. Our numerics thus shows that these neglected higher-order terms can be absorbed in a renormalization of the Luttinger parameter $K$. For quantitative predictions in a potential experimental realization, one could similarly extract $K$ (which determines the coupling $\beta$ in the sG model through $\beta^{2}=\pi\kappa^{2}K$) by a fit of Eq. (III.2.1) instead of relying on the above theoretical estimates. We also emphasize that larger values of $\beta^{2}$ become accessible by realizing interactions with $\kappa>1$. In App. E we further investigate the impact of next-nearest-neighbors (NNN) terms, and show that they lead to the same Luttinger liquid model, yet with reduced deviations in the Luttinger parameter $K$.
@@ III.2.2 The sine-Gordon model
which we test in our numerics by a linear fit in log-log scale. More precisely, we fit data points corresponding to a given $\beta^{2}$, while we vary over $M_{0}^{\prime}$, thus resulting in different values of the observables. The data points considered are obtained by the double extrapolation in $J$ and $N$. The result of the linear fit is shown in Fig. 4(a), where we compare the fit results with the scaling dimensions extracted from the CFT. The observed agreement corroborates our identification of the renormalized coupling $\beta^{2}$. In App. F.1, we further investigate the impact of next-nearest-neighbor interaction terms on the value of $\beta^{2}$, and show that they lead to a reduced renormalization of the coupling, as already observed for the Luttinger parameter $K$.
Since the sG model has no further coupling parameters, we are left with the renormalization scale, which is set by the bare mass $M_{0}^{\prime}$ on the microscopic lattice. Given the observed renormalization of $\beta^{2}$ due to higher derivatives from the term $\propto\cos\left[\beta\left(\hat{\varphi}_{i}-\hat{\varphi}_{j}\right)/\kappa\right]$ [see App. F.2], we depart from the microscopic identification given in Eq. (9). Instead, we fix $M_{0}^{\prime}$ through the formally equivalent relation
@@
which is, however, independent of the microscopic parameter $V^{\prime}_{\text{nn}}$. We obtain an adjusted prediction of $M_{0}^{\prime}$ by inserting the value of $\beta^{\prime}=\beta$ from the previously extracted $\beta^{2}$ [either through $\beta^{2}=\pi\kappa^{2}K$ or Eq. (20)]. As shown in Fig. 2(c), the numerically obtained mass gap $\Delta E$ as a function of mass scale set by this procedure shows excellent agreement with the analytical predictions from Eqs. (12) and (13). This clearly indicates that our numerical results have reached the scaling limit and indeed realize the sG QFT with correctly extracted parameters $M_{0}^{\prime}$ and $\beta^{2}$. Further details about the matching of the mass scale can be found in App. F.3.
Figure 4: Renormalized sG QFT. (a) Coupling $\beta^{2}$ as a function of the microscopic parameters $\kappa^{2}\sqrt{\chi/V_{\text{nn}}^{\prime}}$. The colored dots resp. the yellow diamonds represent the $\beta^{2}$ values extracted in the sG model for $\kappa=1$ resp. $\kappa=2$ with a fit according to Eq. (20) [see inset]. The extracted coupling parameters agree with the CFT relation $\beta^{2}=\pi\kappa^{2}K$ (gray squares for $\kappa=1$, and black triangles for $\kappa=2$), but show deviations from the naive estimate from the microscopic parameters [Eq. (9), dashed black line], signalling a renormalization of $\beta^{2}$. (b) Ground-state energy density $E_{0,\text{latt}}$ (in lattice units) measured with respect to the CFT. For the same value of $\beta^{2}$ we display the energy gap $\Delta E_{\text{sG}}$ (in sG units) in the inset. (c) Expectation value of the vertex operator $\langle e^{i\beta\hat{\varphi}}\rangle$ at $\kappa=1$ versus $M_{0}^{\prime}$. Numerical results (colored plus) and corresponding theoretical predictions (dashed black line). As discussed in the main text, the bare numerical result is renormalized with a factor $Z_{\beta}\approx 1+\mathcal{O}(\beta^{2})$ [see inset]. (d) The expectation value of higher powers of the vertex operators $\langle e^{in\beta\hat{\varphi}}\rangle$ (dark plus) at $\lambda^{\prime}_{\kappa}/V_{\text{nn}}^{\prime}=0.0025$ agree with the Lukyanov-Zamolodchikov conjecture (green circles; dashed lines are guides to the eye) upon rescaling the numerical results with the factor $(Z_{\beta})^{n^{2}}$ [see main text and App. F.4 for more details].
@@
So far, we have compared our numerics to exact results of the sG QFT. Although the model is integrable, not all quantities are known exactly. In particular, there is an unproven conjecture about the expectation values of $\langle e^{in\beta\hat{\varphi}}\rangle$ for generic $n$ [see App. F.4 for the explicit formula]. Here, we test this conjecture in our simulations for integer $n=1,\dots,6$ [see Fig. 4(d)]. We find quantitative agreement with the conjectured functional form when taking into account a multiplicative factor $Z_{\beta}^{(n)}=\left(Z_{\beta}\right)^{n^{2}}$, which relates the higher powers of the vertex operators to their corresponding lattice regularization $\propto\left(\hat{J}_{+}\right)^{n}$.
Altogether, our findings confirm the expected simulability of the sG model regularized with large spins on a finite lattice, although we are restricted to a coupling region close to the classical limit $\beta^{2}\rightarrow 0$. Similarly to the renormalization of the parameter $\beta^{2}$, we attribute this restriction to undesired higher-derivatives terms in the Taylor expansion of the term $\cos[\beta(\hat{\varphi}_{i}-\hat{\varphi}_{j})/\kappa]$ [see App. F.2]. While the accessible couplings for $\kappa=1$ are limited to the massive regime with $\beta^{2}\ll 4\pi$, the comparison with our results for $\kappa=2$ [see Fig. 4(a)] also illustrates that the accessible region can be systematically enlarged by increasing $\kappa$. Specifically, according to Eq. (9), for $\kappa=5$ we expect to reach the sG model with large $\beta^{2}$ up to $\beta^{2}\gtrsim 8\pi$. While we have not investigated this regime classically because the extrapolations for larger $\kappa\gg 1$ require simulations at increasingly large spin length $J$, an experimental implementation could nevertheless probe this regime.